OpenStudy (anonymous):

Can anyone help with algebra 1

OpenStudy (anonymous):

heres the question Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3. (4 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the integer values of x only between -3 and 3. (4 points) Part C: How can you solve the equation 2-x = 4x + 3 graphically? (2 points)

OpenStudy (anonymous):

Anyone know how to solve?

OpenStudy (welshfella):

both 2-x and 4x - 3 are equal to y therefore 2 - x = 4x - 3 (by The Transitive Property of Equality.)

OpenStudy (anonymous):

Thats for A?

OpenStudy (anonymous):

Once this is done i will give you a medal, but in new so i need to learn how to give one

OpenStudy (welshfella):

thats A done yes

OpenStudy (welshfella):

Oh thats not all the answer. The point of intersection is the solution because at that point the values of x and y satisfy both equations

OpenStudy (welshfella):

table will look like